how much of an alloy that is 30% copper should be mixed with 200 ounces of an alloy that is 60% copper in order to get an alloy that is 40% copper?
Answer by mananth(16946)

--- percent ---------------- Amount
Copper Alloy I 60 ---------------- 200 Oz
copper qlloy II 30 ---------------- x Oz
Mixture 40 ---------------- 200 + x Oz

60 * 200 + 30 x = 40 ( 200 + x )
12000 + 30 x = 8000 + 40 x
30 x -40 x = -12000 + 8000
-10 x = -4000
/ -10
x = 400 Oz copper qlloy II

m.ananth@hotmail.ca